OpenStudy (anonymous):

Solve the inequalities. Express your answer in interval notation. x^2-8x+12>0 Solve the inequalities. Express your answer in interval notation. x^2-8x+12>0 @Mathematics

OpenStudy (anonymous):

\[ x<2\text{ or }x>6 \]

OpenStudy (mathmagician):

x 2 −8x+12=0 D=64−48=16 x 1,2 =(8±4)/2=6;2 when x<2, or x>6 x 2 −8x+12>0 , so the interval is (2;6)

OpenStudy (mathmagician):

sorry again, \[(-\infty;2) and(6;\infty)\]

OpenStudy (anonymous):

\[(-\infty,2)\cup(6,\infty)\]